OpenStudy (anonymous):

help please ! find the GCF of 12^a b^5 c^4 and 30a^3 b^5 c

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

is that 12^a written correctly?

OpenStudy (anonymous):

oh no sorry its 12a^2

OpenStudy (cwrw238):

can you figure out the GCF of 12 and 30?

OpenStudy (cwrw238):

the GCF is the highest number which will divide exactly into 12 and 30

OpenStudy (anonymous):

so what wil the gcf be ?

OpenStudy (cwrw238):

a good way to find these is to factor each number 12 = 2*2*3 30 = 2*3*5 note that 2 and 3 match up in both factors so gcf = 2*3 = 6 similarly the GCF of a^2 and a^3 = a^2 and of b^5 and b^5 = b^5 what about c^4 and c?